OpenStudy (anonymous):

xe^x+e^x=0 solve for x show me the steps please

OpenStudy (anonymous):

e^x (x + 1) = 0 x = -1

OpenStudy (anonymous):

start with factoring \[e^x(x+1)=0\]and then note that \[e^x>0\] so only \[x+1=0,x=-1\]